OpenStudy (anonymous):

intergrate sin(x)^(1/2) * x^(1/2)

OpenStudy (anonymous):

Is it \[\int\limits_{}^{}\sin (\sqrt{x})\sqrt x dx?\]

OpenStudy (anonymous):

yes. =D

OpenStudy (anonymous):

Start by using a substitution u=x^(1/2), you will get: \[u=\sqrt x \implies du={1 \over 2\sqrt x}dx\], then the integral:\[\int\limits_{}^{}2u^2 \sin {u} {du}\] This can be integrated using integration by parts.